At the beginning of the year, Morales Company had total assets of $833,000 and total liabilities of $553,000. (Treat each item independently.)

(a) If total assets increased $147,000 during the year and total liabilities decreased $82,000, what is the amount of stockholders’ equity at the end of the year?

Stockholders’ equity $enter a dollar amount



(b) During the year, total liabilities increased $105,000 and stockholders’ equity decreased $76,000. What is the amount of total assets at the end of the year?

Total assets $enter a dollar amount



(c) If total assets decreased $88,000 and stockholders’ equity increased $118,000 during the year, what is the amount of total liabilities at the end of the year?

Total liabilities $enter a dollar amount
Answer #1

Total assets=Total liabilities+Total equity

Total equity at beginning=(833,000-553,000)=$280,000

a.Total assets at end=(833,000+147000)=$980,000

Total liabilities would be=(553,000-82000)=$471,000

Hence total equity would be=(980,000-471000)=$509,000

b.

Total equity would be=(280,000-76000)=$204000

Total liabilities would be=(553000+105000)=$658000

Hence Total assets would be=(204000+658000)=$862,000

c.

Total assets would be=(833,000-88000)=$745,000

Total equity would be=(280,000+118,000)=$398000

Hence total liabilities would be=$745,000-$398000

=$347,000